need help on both 26 and 27

26 i need help
A 25-kg box rests on a horizontal surface. the coefticient or stac ition betwweetthe Dk and the sue applied to the box, but the box does not move. What is the friction force from the surface that is exerted on the box? Record Answer Question 27 (1 point) A wild young driver is traveling at a constant speed of 56 m/s drives in a circle track that is 500 m across. What is the size of the acceleration of the car?
Question 26 (1 point) A 25-kg box rests on a horizontal surface. The coefficient of static friction between the box and the surface is 0.40. A horizontal 72-N force is applied to the box, but the box does not move. What is the friction force from the surface that is exerted on the box? Record Answer Question 27 (t point) d young driver is traveling at a constant speed of 59 m/s drives in a circle track that is 5o0 m across What is the size ofthe acceleraton of the car?
Answer #1

Ques 26:

If the box does not move even after applying 72 N horizontal force on it, the force of static friction from the surface must be equal to the horizontal force applied on the box and is equal to 72N (ans)

Ques 27:

Diameter of the circular path, d =500m

radius of the circular path, r = (d/2) =250 m

speed of the car, v = 56 m/s

magnitude of acceleration, a = (v2 / r) = ((56)2 / 250) = 12.544 m/s2 (ans)